# One Token Can Help! Learning Scalable and Pluggable Virtual Tokens for Retrieval-Augmented Large Language Models

### v3

@@ 4.2 Baseline Methods
$\bullet$ Prompt: This method uses a manually-crafted prompt to indicate the use of retrieval information (details are provided in Appendix C).
$\bullet$ LoRA Hu et al. (2022): This method uses LoRA to fine-tune the backbone models. We use the hyperparameters suggested by the LLaMA’s official guidance. [^3] LLaMA Recipes, https://github.com/meta-llama/llama-recipes/blob/main/src/llama_recipes/configs/peft.py
$\bullet$ Prefix-tuning Li and Liang (2021): This method uses prefix-tuning to fine-tune the backbone models. To make a fair comparison with our method, we add 50 prefix tokens for training.
The implementation details are provided in Appendix D.
4.3 Experimental Results
We fine-tune the prefix-tuning, LoRA, and SPRING methods on RAG tasks, and then evaluate their performance in scenarios both with (RAG) and without (non-RAG) retrieval. For SPRING, we use $k=50$ virtual tokens for inference by default, and the impact of token quantity $k$ is discussed in Section 4.4.2. The experimental results are shown in Table 1. To save space, we only show the results based on Mistral-7b-instruct, and other results are provided in Appendix E.
We can observe: (1) It is evident that SPRING significantly improves the RAG performance of the original LLM with manually-crafted prompts (the average EM and F1 scores are improved by $33.9\%$ and $12.7\%$, respectively). It even outperforms LoRA on certain datasets, such as TriviaQA and CoQA. Given that SPRING involves only 0.2M trainable parameters, these results demonstrate its remarkable efficiency and effectiveness. (2) While LoRA achieves slightly better performance, it adjusts the LLMs’ original parameters, which adversely impact their performance in non-RAG scenarios — a significant drop has been observed, far worse than the original models. This challenge also extends to other general generation tasks, which will be discussed in the next section. (3) Using manually-crafted prompts is effective for improving LLMs’ performance on RAG tasks. However, this improvement is limited as no training is involved. (4) All backbone models show improvements with SPRING (see Appendix E), demonstrating its adaptability across various backbone models. (5) SPRING achieves the best performance on the held-out test set, PopQA, validating the good generalizability of our method. (6) Interestingly, prefix-tuning cannot perform well for RAG, highlighting that the insertion position of the virtual tokens in SPRING is both reasonable and effective. We conduct more analysis in Appendix F.
4.4 Further Analysis
@@
4.4.1 Performance on Other Tasks
To examine the impact of different fine-tuning methods on the inherent capabilities of LLMs, we evaluate the performance of models fine-tuned by LoRA and SPRING on several other (non-RAG) tasks. These tasks are commonly used to evaluate LLMs’ reasoning, mathematical abilities, and world knowledge, including BoolQ Clark et al. (2019), CommonsenseQA Talmor et al. (2019), GSM8K Cobbe et al. (2021), and MMLU Hendrycks et al. (2021). The experimental results are shown in Table 2.444We notice that our results are quite different from those officially reported, which we attribute to the impact of different prompts. Since official prompts for testing are unavailable, we provide the prompts we used in Appendix C to facilitate reproducibility. From the results, we can observe: (1) Thanks to the plug-and-play design of our method, SPRING can revert to to the original LLMs by not using virtual tokens. Therefore, it successfully preserves the original capabilities of the LLMs. In contrast, LoRA, which adjusts the model’s parameters for RAG tasks, inevitably compromises the model’s performance on other tasks. (2) A noticeable decline is observed in the few-shot evaluation, reflecting a decrease in the in-context learning abilities of LLMs. This decline may stem from the fact that RAG fine-tuning does not incorporate in-context learning capabilities. Besides, fine-tuning for RAG tasks may lead the model to overfit to specific task formats, thereby impairing its general generation abilities (more empirical studies are detailed in Appendix G).
@@ 4.4.1 Performance on Other Tasks
4.4.2 Impact of Token Quantity
In SPRING, we design a scalable training approach that enables to use arbitrary numbers of virtual tokens in inference. To validate its effectiveness, we test the performance of our method with various numbers of virtual tokens and compare it with a variant model trained with a fixed number of tokens ($k=50$). The experimental results are illustrated in Figure 4. In general, we can observe that the performance of SPRING increases with more virtual tokens used. Surprisingly, SPRING can significantly enhance LLMs’ performance in RAG scenarios with just a single token, which is very encouraging. [^5] This varies across different LLMs (see Appendix H). In comparison, training with a fixed number of tokens limits the flexibility of SPRING, making it can only be used with the same number of tokens in inference (i.e., $k=50$).
@@ 4.4.2 Impact of Token Quantity
4.4.3 Effects of Different Retrievers
@@
4.4.4 Influence of Retrieved Passages
During the fine-tuning of SPRING, we construct training samples by randomly selecting the top-$m$ ($m\in[1,5]$) retrieved passages. This aims to enhance SPRING’s adaptability by ensuring it can operate effectively with varying numbers of retrieved passages in real-world scenarios. To evaluate the effect of this training strategy, we test the SPRING’s performance across a range from zero to five passages. Figure 5 illustrates the results. We can find that SPRING’s performance gradually improves as more retrieved passages are used ($m=0\rightarrow 4$), suggesting that more retrieved passages contribute valuable knowledge for question answering. However, the performance peaks at four passages and declines when more passages are added. This decrease could be attributed to noise accumulation within the retrieved knowledge, a phenomenon also reported in recent studies Yoran et al. (2023); Wu et al. (2024). Despite this, the use of retrieved passages still results in performance gains compared to scenarios without retrieval ($m=0$), highlighting again the benefits of RAG.
@@ 4.4.4 Influence of Retrieved Passages
4.4.5 Cross-Dataset Generalizability
Inspired by previous studies in multi-task learning Raffel et al. (2020); Khashabi et al. (2020), we mix eight QA datasets for training as they require similar LLM capabilities (e.g., reasoning). To study the impact of this strategy, we conduct experiments by training SPRING on each dataset individually and then testing its performance on the others. Table 4 shows partial results, and more results are available at Appendix E. As indicated, training on a mixed dataset generally enhances performance on most datasets, thereby validating the benefits of multi-task learning. While training on a single dataset, such as NQ, may yield superior results on its specific test set, such improvements often fail to generalize to other datasets. Notably, training solely on NQ may negatively impact performance on MS MARCO, where the original LLM using a prompt could outperform it. These findings inspire us to carefully consider the interaction between different datasets when applying our method in future applications.
5 Conclusion
@@
Limitations
In this work, we introduced a new parameter-efficient fine-tuning method for RAG. By employing scalable and pluggable virtual tokens, our method not only significantly enhances the RAG performance of LLMs but also preserves their general capabilities. However, there are still several limitations in our current work that offer avenues for future research.
First, as discussed in Section 4.4.2, while increasing the number of virtual tokens generally leads to improvements, these benefits are subject to the law of diminishing marginal utility, indicating a performance upper-bound for our method. The potential reason is that our SPRING only fine-tunes the embeddings of the added tokens, which may have limited influence on LLMs’ inference process. Second, compared to LoRA, which optimizes LLMs’ parameters in each layer of LLMs, our SPRING performs slightly worse. This is potentially because our optimization is restricted to the embedding layer only. Therefore, it is interesting to explore a method that optimizes more parameters of LLMs but can still work in a plug-and-play manner. Finally, due to the limited computational resources, our main experiments are conducted on models up to 7B parameters. Although we have demonstrated the effectiveness of our method on 13B models, its performance on larger models, such as those with 30B or 70B parameters, remains untested but holds promise.
Appendix A Example Code for Usage
We provide an example of how to use our SPRING in practice. During the initialization phase, SPRING merges the trained embeddings with the original embeddings of the LLM, and assigns these new embeddings to newly added special tokens. These special tokens are then inserted between the retrieved results and the user input for generation.
QA Dataset # Training # Test Average Input Length Average Output Length
Trivia QA 78,785 11,313 20.93 5.16
Natural Questions 79,168 3,610 11.31 4.66
Hotpot QA 90,447 7,405 26.13 4.38
SQuAD 1.1 87,599 10,570 13.89 5.56
Web Questions 3,778 2,032 10.04 4.10
2WikiMultiHopQA 15,000 12,576 19.33 4.10
CoQA 7,199 500 7.89 3.89
MS MARCO 80,000 55,636 7.76 20.75
PopQA - 14,267 9.24 3.48
BoolQ - 3,270 10.70 1.00
CommonsenseQA - 1,221 15.91 1.00
GSM8K - 1,319 64.90 3.29
MMLU - 14,042 63.91 9.62
Retrieval Set # Passages Average Length
Wikipedia Passages 21,015,324 169.57
MS MARCO Passages 8,841,823 95.73
Table 5: Statistics of all datsaets. The input/output length is the number of tokens calculated by Mistral tokenizer.
Appendix B Datasets
We use the following datasets in our experiments, and their statistics are shown in Table 5.
(1) TriviaQA Joshi et al. (2017) is a reading comprehension dataset including question-answering pairs created by trivia enthusiasts, accompanied by independently gathered evidence documents. The questions are relatively complex and compositional, which require advanced reasoning to derive answers. This dataset is licensed under Apache 2.0 License.
(2) Natural Questions Kwiatkowski et al. (2019) is a question-answering dataset. The questions are collected from the Google search engine, with answers annotated based on the Wikipedia pages. All questions are from real users. This dataset is licensed under Apache 2.0 License.
(3) HotpotQA Yang et al. (2018) is a question-answering dataset with natural and multi-hop questions. All question-answer pairs are Wikipedia-based, where the questions are diverse and include factoid comparison questions. This dataset is licensed under CC BY-SA 4.0 License.
(4) SQuAD 1.1 Rajpurkar et al. (2016) is a reading comprehension dataset derived from Wikipedia articles. The questions are posed by crowd workers while the answers are text segments from the corresponding passages. This dataset is licensed under CC BY-SA 4.0 License.
(5) Web Questions Berant et al. (2013) is a question-answering dataset using Freebase as the knowledge base. The questions are collected via the Google Suggest API, and the answers are defined as Freebase entities. This dataset is licensed under CC BY 4.0 License.
(6) 2WikiMultiHopQA Ho et al. (2020) is a multi-hop question-answering dataset designed to test reasoning and inference skills. It introduces evidence information containing a reasoning path to ensure the quality and multi-hop nature of the questions. This dataset is licensed under Apache 2.0 License.
(7) CoQA Reddy et al. (2019) is a conversational question-answering dataset, which is obtained from human conversations about text passages from various domains. To keep the completeness of the question, we only use the first-round questions and their corresponding answers. This dataset contains data from several sources, under the license of CC BY-SA 4.0, MSR-LA, Apache, and RACE-specific License.
(8) MS MARCO Nguyen et al. (2016) is a question-answering dataset featuring real Bing questions and human-generated answers. Since MS MARCO is extremely large, to balance the data amount in each dataset, we randomly sample 80,000 examples from the training set for fine-tuning. Besides, we also filter out questions without answers in the test set for evaluation. This dataset is licensed under the MIT License.
(9) PopQA Mallen et al. (2023) is an open-domain QA dataset with fine-grained and long-tail Wikidata entities, Wikipedia page views, and the information of relationship type. Since LLMs often struggle with less popular factual knowledge, retrieval augmentation is particularly beneficial for this dataset. This dataset is licensed under the MIT License.
(10) BoolQ Clark et al. (2019) is a reading comprehension dataset, where each sample consists of a question, a passage excerpt, and a yes/no answer. The questions are often complex and require difficult entailment-like inference to solve. Therefore, it is often used to evaluate LLMs’ text comprehension capability. This dataset is licensed under CC BY-SA 3.0 License.
(11) CommonsenseQA Talmor et al. (2019) is a commonsense question-answering dataset with multiple-choice questions designed to distinguish between related concepts. It is commonly used for evaluating LLMs’ commonsense reasoning ability. This dataset is licensed under the MIT License.
(12) GSM8K Cobbe et al. (2021) is a dataset that contains high-quality linguistically diverse grade school math word problems to measure models’ ability of multi-step mathematical reasoning. This dataset is licensed under the MIT License.
(13) MMLU Hendrycks et al. (2021) is a comprehensive aggregated benchmark. It contains 57 tasks to measure LLMs’ knowledge across different subjects. This dataset is licensed under the MIT License.
Figure 7: The prompt used in the baseline methods. “{x}” is the placeholder which will be replaced by the real data.
Figure 8: The prompt used in the BoolQ and CommonsenseQA datasets. “{x}” is the placeholder which will be replaced by the real data.
Figure 9: The prompt used in the GSM8K and MMLU datasets. “{x}” is the placeholder which will be replaced by the real data.
Appendix C Prompt
Figure 7 illustrates the prompts used in the baseline methods. In the Concat method, we simply concatenate the retrieval results and the question for evaluation. In the Prompt method, we use a manually-crafted prompt to indicate the incorporation of retrieval information.
Figure 8 and Figure 9 present the prompts used for the BoolQ, CommonsenseQA, GSM8K, and MMLU datasets. We follow the practices in lm-evaluation-harness and adopt different strategies for performing evaluation on these datasets. [^6] https://github.com/EleutherAI/lm-evaluation-harness Specifically, for the BoolQ dataset, we compare the generation probabilities of the token “true” and “false”, and then we select the token with the higher probability as the predicted result. For CommonsenseQA and MMLU, we determine the answer by comparing the generation probabilities of each option with its content and choosing the option with the highest probability. For GSM8K, we provide LLMs with chain-of-thought demonstrations and extract the answer from the generated text. It is worth noting that the choice of prompt significantly affects the performance of LLMs, so we select the most effective prompt for the vanilla model and apply it across other methods for evaluation. We also observe that our results are worse than those officially reported, and we attribute the gap to the different prompts used for evaluation. Unfortunately, the official prompts are not released publicly, so we are limited to selecting the best-performing prompt from our experiments.
Appendix D Implementation Details
[…]
Mistral-7b-base Mistral-7b-instruct
Prompt SPRING Prompt SPRING
EM F1 EM F1 EM F1 EM F1
TQA 0.00 17.03 68.54 86.89 57.79 80.33 65.71 85.26
NQ 0.08 7.47 46.90 73.28 28.99 58.72 42.35 70.73
HQA 0.00 12.22 39.78 68.48 26.36 56.15 35.26 65.44
SQuAD 0.01 12.04 37.13 69.02 23.92 57.66 33.67 66.99
WebQ 0.00 11.61 35.25 66.44 17.53 52.10 31.84 64.78
2Wiki 0.00 14.48 35.42 64.13 22.64 55.58 31.80 62.03
CoQA 0.00 8.15 15.23 43.94 8.20 36.72 13.28 42.41
MS MARCO 0.08 21.21 7.59 53.90 5.73 53.81 6.57 53.48
PopQA 0.00 14.72 50.37 74.84 39.79 68.26 48.71 73.90
Average 0.02 13.21 37.36 66.77 25.66 57.70 34.35 65.00
Table 6: Performance of SPRING with Mistral-7b models.
LLaMA-2-7b-base LLaMA-2-7b-chat
Prompt SPRING Prompt SPRING
EM F1 EM F1 EM F1 EM F1
TQA 0.01 34.37 65.42 85.30 59.84 81.75 66.95 85.59
NQ 0.00 22.31 43.48 71.04 32.46 62.10 42.13 70.57
HQA 0.00 26.17 35.48 65.59 26.39 56.58 33.88 64.47
SQuAD 0.00 26.49 35.10 67.34 24.57 59.40 33.36 66.60
WebQ 0.00 23.76 31.20 63.73 20.85 57.22 31.10 64.31
2Wiki 0.00 25.83 30.61 60.90 20.69 55.82 30.86 61.45
CoQA 0.00 16.08 13.67 42.73 8.59 38.56 14.06 43.99
MS MARCO 0.00 51.75 7.43 54.11 5.32 48.42 7.18 52.85
PopQA 0.00 32.17 49.93 74.41 45.08 72.23 49.00 73.98
Average 0.00 28.77 34.70 65.02 27.09 59.12 34.28 64.87
Table 7: Performance of SPRING with LLaMA-2-7b models.
LLaMA-2-13b-base LLaMA-2-13b-chat
Prompt SPRING Prompt SPRING
EM F1 EM F1 EM F1 EM F1
TQA 0.26 30.77 66.95 86.48 63.54 82.54 68.52 86.62
NQ 0.57 19.54 44.91 71.70 36.66 67.03 43.80 71.15
HQA 0.07 19.72 37.24 66.81 29.27 59.36 36.23 66.11
SQuAD 0.26 22.79 36.35 68.48 28.45 61.92 34.80 67.76
WebQ 0.24 20.77 32.32 64.54 20.61 55.75 31.01 64.17
2Wiki 0.10 22.63 33.32 62.85 28.01 59.20 32.32 62.43
CoQA 0.39 14.85 14.84 44.00 9.96 39.27 14.26 44.49
MS MARCO 0.57 47.30 7.36 53.71 6.24 47.68 7.51 52.94
PopQA 0.13 24.06 50.08 74.72 42.49 69.67 50.29 74.84
Average 0.29 24.71 35.93 65.92 29.47 60.27 35.42 65.61
Table 8: Performance of SPRING with LLaMA-2-13b models.
Training $\rightarrow$ TQA NQ HQA SQuAD Mix
Test $\downarrow$ EM F1 EM F1 EM F1 EM F1 EM F1
TQA 62.80 84.65 65.51 84.73 66.42 85.26 67.48 85.76 65.71 85.26
NQ 32.19 64.98 45.26 72.68 40.79 70.18 37.88 67.23 42.35 70.73
HQA 22.97 56.95 28.11 59.45 36.30 66.33 30.43 61.46 35.26 65.44
SQuAD 25.24 61.39 30.81 64.32 33.37 66.36 35.01 67.83 33.67 66.99
WebQ 26.03 62.40 34.13 67.30 29.15 64.04 28.27 63.38 31.84 64.78
2Wiki 17.32 54.41 25.43 58.21 30.42 61.62 25.48 58.26 31.80 62.03
CoQA 6.05 36.61 7.62 38.54 12.50 43.12 7.62 37.74 13.28 42.41
MS MARCO 3.26 31.88 3.88 33.33 5.18 36.12 3.89 45.80 6.57 53.48
PopQA 44.07 72.50 48.28 74.07 48.80 74.32 48.40 73.83 48.71 73.90
Average 26.66 58.42 32.11 61.40 33.66 63.04 31.61 62.37 34.35 65.00
Table 9: Full results of training on a specific dataset and a mixture of all datasets.
Appendix E Full Experimental Results
In the main body of our paper, we present the results using Mistral-7b-instruct as the backbone model. We also apply our method to Mistral-7b-base, LLaMA-2-7b-chat, LLaMA-2-7b-base, LLaMA-2-13b-base, and LLaMA-2-13b-chat. Their results are shown in Table 6, Table 7, and Table 8. We can observe that SPRING can achieve consistent improvement for all models. Specifically, using prompt is very effective for instruction-tuned models (i.e., Mistral-instruct and LLaMA-chat) but less so for base models. This is because the base model cannot understand the instructions to read the retrieved results and provide correct answers. Mistral series models perform slightly better than LLaMA series models of the same size. This is consistent with their officially reported benchmark performance. Furthermore, larger models tend to perform better than smaller models, and our SPRING can bring improvement across models of different sizes. In future, we plan to extend our method to much larger models (such as 30B or 70B) when more computing resources available.
$[T,R,Q]$ SPRING $[R,Q,T]$
EM F1 EM F1 EM F1
TQA 11.74 59.97 65.71 85.26 64.90 84.65
NQ 13.04 38.22 42.35 70.73 43.16 71.30
HQA 5.79 42.59 35.26 65.44 34.00 64.27
SQuAD 7.19 39.75 33.67 66.99 34.09 66.75
WebQ 4.44 31.55 31.84 64.78 31.35 64.84
2Wiki 4.38 41.82 31.80 62.03 30.64 61.01
CoQA 1.56 20.02 13.28 42.41 12.30 41.18
MS MARCO 0.60 50.56 6.57 53.48 6.94 49.40
PopQA 10.02 44.54 48.71 73.90 46.55 72.33
Average 6.53 41.00 34.35 65.00 33.77 63.97
Table 10: Performance of different virtual token insertion positions. The best results are in bold.
Appendix F Impact of Token Insertion Position
In SPRING, virtual tokens ($T$) are strategically inserted between the retrieved results ($R$) and the question ($Q$). Generally, there are three possible insertion positions:
  • $[T,R,Q]$: The virtual tokens are added at the beginning of the sequence, akin to prefix-tuning. In this setup, due to the auto-regressive nature of LLMs, these tokens cannot access subsequent sequence information.
  • $[R,T,Q]$: This is what we used in SPRING. The virtual tokens can attend to the retrieved results, while positioning the question immediately next to the answer.
  • $[R,Q,T]$: The virtual tokens are added at the end of the sequence. They are able to attend to all preceding information, including both the retrieved results and the question.
We compare the performance of these three strategies. The results are shown in Table 10. It is evident that SPRING achieves the best performance among the three strategies, which demonstrates the effectiveness of our token placement design. Placing virtual tokens at the beginning of the sequence (prefix-tuning) is the worst strategy, because the added tokens cannot adapt based on the retrieved results or the question during fine-tuning. This result is different from those reported in the original prefix-tuning study Li and Liang (2021), which suggested that prefix-tuning can yield better performance. We attribute this discrepancy to the different targets of general generation tasks versus RAG tasks. In general generation, all contexts are crucial for predicting the next token, hence adding trainable tokens at the beginning can effectively influence the activation of contexts. In contrast, in RAG tasks, the retrieved information is supplementary to answering the user’s question, so the virtual tokens are added to help LLMs understand and leverage the retrieved information rather than enhance the retrieved information itself. This speculation is further validated by the strong performance achieved by adding tokens at the end of the sequence, where they can also access the retrieved results. Among the three strategies, SPRING is the most effective, not only allowing the virtual tokens to utilize the retrieved information optimally but also maintaining coherence between the question and the answer being generated.
Prompt 1:
[…]
Prompt 2:
[…]
Prompt 3:
[…]
Prompt 4:
Table 11: The prompts we used for exploring their impacts on inference.
Figure 10: Performance under various prompts.
Appendix G Impact of Prompt after Fine-tuning
In our experiments, we notice that models fine-tuned by certain methods (such as LoRA) are very sensitive to the prompts used for inference. To investigate this problem, we conduct an empirical study. We mainly focus on the impact of the prompt $P$ inserted between the retrieved results and the question, formatted as $[R;P;Q]$. We consider four different prompts listed in Table 11. Specifically, for LoRA, we use the first prompt during training, and test its performance with all prompts. For SPRING, it does not need a prompt in training, which can be seen as using the fourth prompt (none). We also evaluate its performance with the four prompts.
The experimental results are shown in Figure 10. It clearly reflects the prompt dependency of models fine-tuned with LoRA. LoRA achieves the highest EM score of $37.13$ with the first prompt, which is used during its training. The performance remains robust with the second prompt, which is very similar to the first but omits the last sentence. However, the third prompt, while semantically similar to the first, involves different phrasing and significantly impacts LoRA’s performance. Notably, in scenarios where no prompt is used (the fourth prompt), LoRA’s performance drops sharply, even worse than that of the original model. These results indicate that fine-tuning with LoRA may lead to overfitting problems. This may also be the reason why models tuned with LoRA cannot perform well on non-RAG tasks. In comparison, our SPRING demonstrates a consistent performance across all prompts, demonstrating its robustness and generalizability.
Appendix H Impact of Token Quantity
In the main body of our paper, we explore the impact of token quantity based on the Mistral-7b-instruct model. Since this impact may vary across different models, we also extend our analysis to include the LLaMA-2-7b-chat model. The result is shown in Figure 11. It can be seen that the general trend is similar to that based on the Mistral model, namely the performance of SPRING is gradually improved as more virtual tokens are used. However, LLaMA requires at least three tokens to obtain improvement over the original model with prompts. This indicates that different models may require various numbers of virtual tokens to effectively learn and perform RAG tasks.

### v4

@@ Introduction
$\bullet$ Prompt: This method uses a manually-crafted prompt to indicate the use of retrieval information (details are provided in Appendix).
$\bullet$ Prefix-tuning (Li and Liang 2021): This method uses prefix-tuning to fine-tune the backbone models. To make a fair comparison with our method, we add 50 prefix tokens for training.
$\bullet$ LoRA (Hu et al. 2022): This method uses LoRA to fine-tune the backbone models. We use the hyperparameters suggested by the LLaMA’s official guidance. [^2] LLaMA Recipes, https://github.com/meta-llama/llama-recipes/blob/main/src/llama˙recipes/configs/peft.py To further validates the effectiveness of our SPRING on models that have already been optimized for RAG, we also train our method based on the LoRA checkpoint, and denote this variant as SPRING+.
Implementation Details
[…]
Experimental Results
We fine-tune the prefix-tuning, LoRA, and SPRING methods on RAG tasks, and then evaluate their performance in scenarios both with (RAG) and without (non-RAG) retrieval. For SPRING, we use $k=50$ virtual tokens for inference by default, and the impact of token quantity $k$ is discussed in later. The experimental results are shown in Table 1. To save space, we only show the results based on Mistral-7b-instruct, and other results are provided in Appendix.
We can observe: (1) It is evident that SPRING significantly improves the RAG performance of the original LLM with manually-crafted prompts (the average EM and F1 scores are improved by $43.4\%$ and $17.3\%$, respectively). It outperforms LoRA on certain datasets, such as TriviaQA and CoQA. Given that SPRING involves only 0.2M trainable parameters, these results demonstrate its remarkable efficiency and effectiveness. (2) While LoRA achieves slightly better performance on some datasets, it adjusts the LLMs’ original parameters, which adversely impact their performance in non-RAG scenarios—a significant drop has been observed, even far worse than the original models. This challenge also extends to other general generation tasks, which will be discussed in the next section. (3) In non-RAG evaluation, only SPRING and SPRING+ demonstrate better performance than the Prompt method. This indicates that even in the absence of retrieved results, adding virtual tokens is still beneficial. We speculate that beyond simply utilizing retrieved results, virtual tokens can help the LLM understand the task goal and format (e.g., the task is question-answering rather than text continuation). (4) Based on the LoRA’s checkpoint, SPRING+ achieves the best performance on most datasets. Additionally, all backbone models show improvements with SPRING (see Appendix). These findings verify the versatility and flexibility of our approach, confirming its suitability for enhancing various LLMs in RAG scenarios. (5) Using manually-crafted prompts is effective for improving LLMs’ performance on RAG tasks. However, this improvement is limited as no training is involved. (6) SPRING achieves robust performance on the held-out datasets, validating the good generalizability of our method. (7) Interestingly, prefix-tuning cannot perform well for RAG, highlighting that the insertion position of the virtual tokens in SPRING is both reasonable and effective. We provide more analysis in Appendix.
Further Analysis
@@
Performance on Other Tasks
To examine the impact of different fine-tuning methods on the inherent capabilities of LLMs, we evaluate the performance of models fine-tuned by LoRA and SPRING on several other (non-RAG) tasks. These tasks are commonly used to evaluate LLMs’ reasoning, mathematical abilities, and world knowledge, including BoolQ (Clark et al. 2019), CommonsenseQA (Talmor et al. 2019), GSM8K (Cobbe et al. 2021), and MMLU (Hendrycks et al. 2021). The experimental results are shown in Table 2.333We notice that our results are quite different from those officially reported, which we attribute to the impact of different prompts. Since official prompts for testing are unavailable, we provide the prompts we used in Appendix to facilitate reproducibility. From the results, we can observe: (1) Thanks to the plug-and-play design of our method, SPRING can revert to to the original LLMs by not using virtual tokens. Therefore, it successfully preserves the original capabilities of the LLMs. In contrast, LoRA, which adjusts the model’s parameters for RAG tasks, inevitably compromises the model’s performance on other tasks. (2) A noticeable decline is observed in the few-shot evaluation, reflecting a decrease in the in-context learning abilities of LLMs. This decline may stem from the fact that RAG fine-tuning does not incorporate in-context learning capabilities. Besides, fine-tuning for RAG tasks may lead the model to overfit to specific task formats (prompts), thereby impairing its general generation abilities (more empirical studies are detailed in Appendix).
@@
Impact of Token Quantity
In SPRING, we design a scalable training approach that enables to use arbitrary numbers of virtual tokens in inference. To validate its effectiveness, we test the performance of our method with various numbers of virtual tokens and compare it with a variant model trained with a fixed number of tokens ($k=50$). The experimental results are illustrated in Figure 3. In general, we can observe that the performance of SPRING increases with more virtual tokens used. Surprisingly, SPRING can significantly enhance LLMs’ performance in RAG scenarios with just a single token, which is very encouraging. [^4] This varies across different LLMs (see Appendix). In comparison, training with a fixed number of tokens limits the flexibility of SPRING, making it can only be used with the same number of tokens in inference (i.e., $k=50$).
@@
Effects of Different Retrievers
@@
Influence of Retrieved Passages
During the fine-tuning of SPRING, we construct training samples by randomly selecting the top-$m$ ($m\in[1,5]$) retrieved passages. This aims to enhance SPRING’s adaptability by ensuring it can operate effectively with varying numbers of retrieved passages in real-world scenarios. To evaluate the effect of this training strategy, we test the SPRING’s performance across a range from zero to five passages. Figure 4 illustrates the results. We can find that SPRING’s performance gradually improves as more retrieved passages are used ($m=0\rightarrow 4$), suggesting that more retrieved passages contribute valuable knowledge for question answering. However, the performance peaks at four passages and declines when more passages are added. This decrease could be attributed to noise accumulation within the retrieved knowledge, a phenomenon also reported in recent studies (Yoran et al. 2023; Wu et al. 2024). Despite this, the use of retrieved passages still results in performance gains compared to scenarios without retrieval ($m=0$), highlighting again the benefits of RAG.
@@
Cross-Dataset Generalizability
Inspired by previous studies in multi-task learning (Raffel et al. 2020; Khashabi et al. 2020), we mix eight QA datasets for training as they require similar LLM capabilities (e.g., reasoning). To study the impact of this strategy, we conduct experiments by training SPRING on each dataset individually and then testing its performance on the others. Table 4 shows partial results, and more results are available at Appendix. As indicated, training on a mixed dataset generally enhances performance on most datasets, thereby validating the benefits of multi-task learning. While training on a single dataset, such as NQ, may yield superior results on its specific test set, such improvements often fail to generalize to other datasets. Notably, training solely on NQ may negatively impact performance on MS MARCO, where the original LLM using a prompt could outperform it. These findings inspire us to carefully consider the interaction between different datasets when applying our method in future applications.
Conclusion
@@
Acknowledgment
This work was supported by National Natural Science Foundation of China (Grant No. 62402497 and 62272467), Beijing Natural Science Foundation L233008, and Beijing Municipal Science and Technology Project No. Z231100010323009. The work was partially done at the Engineering Research Center of Next-Generation Intelligent Search and Recommendation, MOE.